# Fine-Tuning Visual Autoregressive Models for Subject-Driven Generation

### v1

@@ Abstract
†† ∗ Corresponding authorRecent advances in text-to-image generative models have enabled numerous practical applications, including subject-driven generation, which fine-tunes pretrained models to capture subject semantics from only a few examples. While diffusion-based models produce high-quality images, their extensive denoising steps result in significant computational overhead, limiting real-world applicability. Visual autoregressive (VAR) models, which predict next-scale tokens rather than spatially adjacent ones, offer significantly faster inference suitable for practical deployment. In this paper, we propose the first VAR-based approach for subject-driven generation. However, naïve fine-tuning VAR leads to computational overhead, language drift, and reduced diversity. To address these challenges, we introduce selective layer tuning to reduce complexity and prior distillation to mitigate language drift. Additionally, we found that the early stages have a greater influence on the generation of subject than the latter stages, which merely synthesize local details. Based on this finding, we propose scale-wise weighted tuning, which prioritizes coarser resolutions for promoting the model to focus on the subject-relevant information instead of local details. Extensive experiments validate that our method significantly outperforms diffusion-based baselines across various metrics and demonstrates its practical usage. Codes are available at github.com/jiwoogit/ARBooth.
@@ 1 Introduction
To effectively address these issues, we introduce selective layer tuning and prior distillation. Through our analysis, we observe that cross-attention and feed-forward layers in the VAR transformer exhibit larger weight updates. As larger updates imply the significance of weights to capture the subject-relevant information, we selectively fine-tune the module mostly updated. This approach substantially reduces computational overhead without sacrificing performance. Additionally, our proposed prior distillation leverages the pretrained model’s semantic knowledge to produce diverse instances of the same subject class without requiring extensive external datasets. By utilizing the efficient inference of the VAR architecture, our approach enables end-to-end fine-tuning, effectively mitigating language drift and improving output diversity.
Furthermore, we analyze the generation impact of different scales in the VAR architecture. Specifically, we observe that coarser-scale token maps have a greater influence on synthesizing the identity of subject, whereas finer-scale tokens have a relatively smaller impact. Motivated by this observation, we propose scale-wise weighted tuning, which prioritizes coarser resolutions to effectively capture subject semantics while placing less emphasis on finer-scale details such as local texture. Extensive experiments demonstrate that our method accurately encodes the target concept while maintaining robust editing capabilities, significantly outperforming diffusion-based baselines.
@@ 4.1 Selective Layer Tuning
To address this, we aim to identify a compact yet effective subset of parameters for fine-tuning, inspired by CustomDiffusion [23]. Specifically, we analyze layer-wise adaptation during personalization by measuring weight changes between the original and fully fine-tuned VAR models. As reported in Fig. 3, feed-forward layers exhibit the most significant weight changes. This suggests that these layers primarily integrate subject-specific information. We first attempt to fine-tune only the feed-forward layers while keeping other layers fixed. However, we empirically observe a deterioration in text alignment, as shown in Fig. 7. To ensure alignment with the target text and subject, we therefore selectively fine-tune only these two components, effectively capturing the subject concept while preserving overall generative capability.
@@ 4.3 Prior Distillation
In our autoregressive framework, we propose an end-to-end prior preserving approach called prior distillation. Unlike diffusion-based methods that necessitate data collection, our approach leverages the fast inference capability of VAR to generate diverse instances of the subject’s class. Specifically, the efficiency of VAR’s inference allow our distillation method to mitigate language drift without the burden of gathering large same-class datasets
@@
In summary, we fine-tune the CA and FFN layers of the transformer while jointly optimizing the scale-wise weighted cross-entropy loss () and prior distillation loss () to enhance subject fidelity and preserve generative consistency.
@@ 5.3 Quantitative Comparison
As shown in Tab. 1, our method outperforms diffusion-based baselines on both Idino and Iclip, indicating improved subject alignment. Notably, it also achieves better performance on Tclip, reflecting enhanced alignment with the text prompt. Although these metrics often involve a trade-off, our method achieves higher results in both, demonstrating a strong enhancement.
In addition, we measure the inference time for generating a single image from a target prompt on an A6000 GPU, reported as the “Time” column in Tab. 1. Our method completes inference in 0.5 seconds, whereas diffusion-based baselines take significantly longer. This speed advantage stems from a shorter progressive inference scheme processing from coarse to fine scales. In contrast, diffusion-based methods require numerous denoising steps within the same dimensional space. These results highlight the practical potential of our approach for real-world applications.
Furthermore, we compare our method with CD [23] and DM [35], which achieve the highest quantitative performance among the baselines in subject fidelity and text alignment, respectively. We conduct a user study with 20 participants, collecting a total of 640 responses from 32 comparative questions. Samples are randomly selected from the experimental dataset. Each question presents real subject images, a text prompt, and three generated images from different methods. Users are asked to select the image that best preserves subject identity and prompt fidelity. As shown in Tab. 2, our method is preferred by (67%, 79%) of users over CD and DM, respectively.
@@ 5.5 Ablation Study
In addition, we compute a prior preservation metric (PRES) [42], defined as the average pairwise similarity between the DINO embeddings of generated images from random subjects of the prior class and real images of our target subject. A higher PRES value indicates that the random subjects resemble the target subject, suggesting that the model’s ability to maintain the original has broken down. Furthermore, we evaluate diversity (DIV) [42] using the average pairwise LPIPS [67] distance between generated images of the same subject with the same prompt. These metrics capture a crucial property by indicating whether the model fails to preserve prior knowledge.
@@
Scale-wise weighted tuning ablation. In our default setting, we use and set the other weight to . To evaluate the impact of weight settings, we ablate using a configuration where all weights are set to . As shown in the third row of Tab. 3 and Fig. 8, scale-wise weighted tuning effectively encourages subject fidelity. Although this results in a slight reduction in text alignment, it also preserves prior knowledge, as evidenced by the PRES metrics. This is mainly because scale-wise tuning emphasizes the more important coarse-scale token maps inherent in visual autoregressive modeling.
@@ 5.6 Additional Analysis
Comparison with baselines in recent diffusion architectures with DreamBooth. Since our base model, Infinity-2B [15], demonstrates strong generative performance compared to Stable Diffusion (SD) 1.4, we ensure a fair comparison by also evaluating DreamBooth [42] on recently released diffusion architectures, including SD 3.5 [1] and FLUX [25], using LoRA [20] plugins. We train all models for the same number of iterations to ours. As shown in Tab. 4 and Fig. 10, our generated images achieve the highest subject fidelity among the baselines. While SD 3.5-medium-LoRA exhibits better text alignment, our approach maintains higher subject fidelity. These results indicate that VAR-based approach could surpass recent diffusion-based baselines with substantially faster inference speed.
Optimization Time. While our method achieves significantly faster inference than diffusion-based baselines, our optimization time is relatively long, taking approximately 6 hours. For reference, DreamBooth [42] applied on SD 1.4 completes optimization in about 30 minutes. However, larger diffusion-based frameworks with comparable performance to Infinity [15], such as SD 3.5 [1] or FLUX [25], require similar training times to ours. We expect that future work exploring efficient encoder-based tuning methods may further reduce end-to-end duration.

### v2

@@ Abstract
†† ∗ Corresponding authorRecent advances in text-to-image generative models have enabled numerous practical applications, including subject-driven generation, which fine-tunes pretrained models to capture subject semantics from only a few examples. While diffusion-based models produce high-quality images, their extensive denoising steps result in significant computational overhead, limiting real-world applicability. Visual autoregressive (VAR) models, which predict next-scale tokens rather than spatially adjacent ones, offer significantly faster inference suitable for practical deployment. In this paper, we propose the first VAR-based approach for subject-driven generation. However, naïve fine-tuning VAR leads to computational overhead, language drift, and reduced diversity. To address these challenges, we introduce selective layer tuning to reduce complexity and prior distillation to mitigate language drift. Additionally, we found that the early stages have a greater influence on the generation of subject than the latter stages, which merely synthesize minor details. Based on this finding, we propose scale-wise weighted tuning, which prioritizes coarser resolutions for promoting the model to focus on the subject-relevant information instead of local details. Extensive experiments validate that our method significantly outperforms diffusion-based baselines across various metrics and demonstrates its practical usage. Codes are available at github.com/jiwoogit/ARBooth.
@@ 1 Introduction
To effectively address these issues, we introduce selective layer tuning and prior distillation. Through our analysis, we observe that cross-attention and feed-forward layers in the VAR transformer exhibit larger weight updates. As larger updates imply the significance of weights to capture the subject-relevant information [35, 26], we selectively fine-tune the module mostly updated. This approach substantially reduces computational overhead without sacrificing performance. Additionally, our proposed prior distillation leverages the pretrained model’s semantic knowledge to produce diverse instances of the same subject class without requiring extensive external datasets. By utilizing the efficient inference of the VAR architecture, our approach enables end-to-end fine-tuning, effectively mitigating language drift and improving output diversity.
Furthermore, we analyze the generation impact of different scales in the VAR architecture. Specifically, we observe that coarser-scale token maps have a greater influence on synthesizing the identity of subject, whereas finer-scale tokens have a relatively smaller impact. Motivated by this observation, we propose scale-wise weighted tuning, which prioritizes coarser resolutions to effectively capture subject semantics while placing less emphasis on minor variations such as finer-scale details. Extensive experiments demonstrate that our method accurately encodes the target concept while maintaining robust editing capabilities, significantly outperforming diffusion-based baselines.
@@ 2 Related Work
Moreover, our analysis also provide valuable insights for future VAR research. For example, we find that FFN layers in VAR are crucial, which aligns closely with recent observation in language models [15, 16]. We expect that these insights can serve as a bridge for transferring techniques from language models to AR-based vision models.
@@ 4.1 Selective Layer Tuning
To address this, we aim to identify a compact yet effective subset of parameters for fine-tuning, inspired by previous works [35, 26]. Specifically, we analyze layer-wise adaptation during personalization by measuring weight changes between the original and fully fine-tuned VAR models. As reported in Fig. 3, feed-forward layers exhibit the most significant weight changes. This suggests that these layers primarily integrate subject-specific information. We first attempt to fine-tune only the feed-forward layers while keeping other layers fixed. However, we empirically observe a deterioration in text alignment, as shown in Fig. 7. To ensure alignment with the target text and subject, we therefore selectively fine-tune only these two components, effectively capturing the subject concept while preserving overall generative capability.
@@ 4.3 Prior Distillation
In our autoregressive framework, we propose an end-to-end prior preserving approach called prior distillation. Unlike diffusion-based methods that necessitate data collection, our approach leverages the fast inference capability of VAR to generate diverse instances of the subject’s class. Specifically, the efficiency of VAR’s inference allow our distillation method to mitigate language drift without the burden of gathering large same-class datasets.
@@
In summary, we fine-tune the CA and FFN layers of the transformer and the text embedding while jointly optimizing the scale-wise weighted cross-entropy loss () and prior distillation loss () to enhance subject fidelity and preserve generative consistency.
@@ 5.2 Evaluation Protocol
For the ablation study, we measure a prior preservation metric (PRES) [48], defined as the average pairwise similarity between the DINO embeddings of generated images from random subjects of the prior class and real images of our target subject. A higher PRES value indicates that the random subjects resemble the target subject, suggesting that the model’s ability to maintain the original has broken down. Furthermore, we compute diversity (DIV) [48] using the average pairwise LPIPS [73] distance between generated images of the same subject with the same prompt. These metrics capture a crucial property by indicating whether the model fails to preserve prior knowledge.
@@ 5.3 Quantitative Comparison
As shown in Tab. 1, our method outperforms diffusion-based baselines on both Idino and Iclip, indicating improved subject alignment. Notably, it also achieves better performance on Tclip, reflecting enhanced alignment with the text prompt. Although these metrics often involve a trade-off, our method achieves higher values in both, demonstrating a strong enhancement.
In addition, we measure the inference time for generating a single image using a target prompt on an A6000 GPU, reported as the “Time” column in Tab. 1. Our method completes the inference in 0.5 seconds, whereas diffusion-based baselines take significantly longer. This speed advantage stems from a shorter progressive inference scheme processing from coarse to fine scales. In contrast, diffusion-based methods require numerous denoising steps within the same dimensional space. These results highlight the practical potential of our approach for real-world applications.
Furthermore, we compare our method with CD [26] and DM [40], which achieve the highest quantitative performance among the baselines in subject fidelity and text alignment, respectively. We conduct a user study with 20 participants, collecting a total of 640 responses from 32 comparative questions. Samples are randomly selected from the experimental dataset. Each question presents real subject images, a text prompt, and three generated images from the baseline methods. Users are asked to select the image that best preserves subject identity and prompt fidelity. As shown in Tab. 2, our method is preferred by (67%, 79%) of users over CD and DM, respectively.
@@ 5.5 Ablation Study
Scale-wise weighted tuning ablation. In our default setting, we use and set the other weight to . To evaluate the impact of weight settings, we ablate using a configuration where all weights are set to . As shown in the third row of Tab. 3 and Fig. 8, scale-wise weighted tuning effectively encourages subject fidelity. Although this results in a slight reduction of text alignment, it also preserves prior knowledge, as evidenced by the PRES metrics. This is mainly because scale-wise weighted tuning emphasizes more important coarse-scale token maps inherent in visual autoregressive modeling.
@@ 5.6 Additional Analysis
Comparison with baselines in recent diffusion architectures with DreamBooth. Since our base model, Infinity-2B [18], demonstrates strong generative performance compared to Stable Diffusion (SD) 1.4, we ensure a fair comparison by also evaluating DreamBooth [48] on recently released diffusion architectures, including SD 3.5 [10] and FLUX [28, 29], using LoRA [23] plugins. We train all models for the same number of iterations as our approach and evaluate on 5 living subject concepts. As shown in Tab. 4 and Fig. 10, our generated images achieve the highest subject fidelity among the baselines. While SD 3.5-medium-LoRA exhibits better text alignment, our approach maintains higher subject fidelity. These results indicate that VAR-based approach could surpass recent diffusion-based baselines with substantially faster inference speed.
Optimization Time. While our method achieves significantly faster inference than diffusion-based baselines, our optimization time is relatively long, taking approximately 6 hours. For reference, DreamBooth [48] applied on SD 1.4 completes optimization in about 30 minutes. However, larger diffusion-based frameworks with comparable performance to Infinity [18], such as SD 3.5 [10] or FLUX [28, 29], require similar training times to ours. We expect that future work exploring efficient encoder-based tuning methods may further reduce end-to-end duration.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by MSIT/IITP (No. RS-2022-II220680, 2020-0-01821, RS-2019-II190421, RS-2024-00459618, RS-2024-00360227, RS-2024-00437633), MSIT/NRF (No. RS-2024-00357729), KNPA/KIPoT (No. RS-2025-25393280), and SEMES-SKKU collaboration funded by SEMES.
Supplementary Material
Appendix A Additional Ablations
A.1 Scale-wise Weighted Tuning for Style Personalization
Unlike diffusion models, where later timesteps capture fine, high-frequency details, our analysis indicates that later scales of large-scale pretrained VAR models contribute minimally to output variation. Consequently, our proposed Scale-wise Weighted Tuning (SWT) strategy de-emphasizes these scales, improving robustness in style personalization tasks, particularly for capturing high-frequency details.
To validate SWT effectiveness, we fine-tune on eight distinct style concepts from DreamBench++ [42], each with nine text prompts, for 100 iterations per style. We generate eight outputs per pair and present representative qualitative comparisons in Fig. 11. As shown in Fig. 11, SWT achieves clearer stylistic expressions, such as distinct swirls and brushstrokes.
A.2 Prior Distillation vs. Prior Preservation Loss
We further compare our Prior Distillation (PD) method against the Prior Preservation Loss (PPL) approach [48]. For integrating PPL into VAR [58], we replace the original MSE objective with cross-entropy objective, using 100 class-specific images generated with the Infinity-2B checkpoint.
Quantitative results in Tab. 5 and qualitative results in Fig. 12 demonstrate comparable performance between PD and PPL across various metrics. Notably, PD eliminates the requirement for same-class dataset collection, substantially reducing the preparation time and computational overhead.
| Variant | Idino | Iclip | Tclip | PRES | DIV |
| Ours w/ PD | 0.786 | 0.853 | 0.267 | 0.709 | 0.272 |
| Ours w/ PPL | 0.653 | 0.822 | 0.265 | 0.608 | 0.378 |
A.3 Selective Layer Tuning with LoRA
We evaluate Selective Layer Tuning (SLT) combined with Prior Distillation (PD) against the popular LoRA method [23]. Specifically, we apply LoRA across all VAR layers at multiple ranks (4 and 16) without employing SLT or PD, to isolate their contributions. As shown in Tab. 6, LoRA without SLT or PD results in unstable fine-tuning, causing visual artifacts or failing to accurately capture subject concepts.
The qualitative results in Fig. 13 further confirm that our combined approach (SLT + PD) stabilizes personalization and preserves subject identity, validating their critical role in achieving robust VAR personalization.
| Base | Adapter | Idino | Iclip | Tclip | PRES | DIV |
| w/o SLT | + LoRA(4) | 0.653 | 0.822 | 0.265 | 0.608 | 0.378 |
| + LoRA(16) | 0.770 | 0.852 | 0.264 | 0.383 | 0.351 | |
| w/o PD | + LoRA(4) | 0.753 | 0.852 | 0.265 | 0.676 | 0.376 |
| + LoRA(16) | 0.775 | 0.854 | 0.265 | 0.781 | 0.373 | |
| Ours (w/ SLT, PD) | 0.786 | 0.853 | 0.267 | 0.709 | 0.272 | |
Appendix B User Study Details
We conduct a user study with 20 participants to evaluate personalization effectiveness. Participants compare outputs from our method against baselines, focusing on subject fidelity and prompt alignment, following the evaluation protocol from DreamBooth [48]. The user interface is shown in Fig. 14.
Appendix C Additional Qualitative Comparisons
C.1 Comparison with FLUX-based Method
We further compare our VAR-based approach with Personalize Anything (PA) [12], which is based on the FLUX framework [29, 28]. As illustrated in Fig. 15, PA [12] struggles to generate dynamic poses from the subject image, whereas our approach successfully synthesizes diverse and dynamic poses, highlighting the efficacy of our VAR-based personalization.
C.2 Comparison with Diffusion-based Methods
Additional qualitative comparisons against diffusion-based baselines are provided in Fig. 16. These examples clearly demonstrate our method’s ability to better preserve subject identity, accurately capturing crucial attributes such as color and shape. Furthermore, our method achieves notably improved alignment with the provided prompts.